40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 6 22:47:41 2021, gautam, Update, ALS, Noisy ALS ALSX_excess.pngbudget.pdf
    Reply  Thu Jan 7 19:16:11 2021, gautam, Update, ALS, Noisy ALS ALS_ERR_MON.pdfAUXnoise.pdf
       Reply  Thu Jan 7 20:07:27 2021, Koji, Update, ALS, Noisy ALS 
          Reply  Thu Jan 7 21:16:22 2021, gautam, Update, ALS, Noisy ALS 
             Reply  Thu Jan 7 23:25:19 2021, Koji, Update, ALS, Noisy ALS 
          Reply  Fri Jan 8 20:01:11 2021, gautam, Update, ALS, Noisy ALS ALSX_IR_green.pdf
Message ID: 15751     Entry time: Wed Jan 6 22:47:41 2021     Reply to this: 15752
Author: gautam 
Type: Update 
Category: ALS 
Subject: Noisy ALS 

Summary:

I want to get back to locking the interferometer so I can test out the newly installed AS WFS. However, the ALS noise is far too high, at least the transition of arm length control from IR to ALS fails reliably with the same settings that worked so reliably previously. I worked on investigating it a bit today.

Timeline

In the latter half of last year, I was focused on the air-BHD setup, so I wasn't checking in on the ALS noise as regularly. 

  1. On Aug 17, the noise was fine.
  2. But on Oct 29, the noise is bad (and it continues to remain so, to the point where I cannot lock the interferometer). 
  3. Koji and Anchal confirmed nothing was touched while they were investigating the ALS system, also on Oct 29. The spectra attached in #15650 don't make any sense to me, the noise at 100 Hz cannot be <100mHz/rtHz. So, inconclusive.

Excess noise:

All tests are done with the arm cavity length locked to the PSL frequency using POX. Then, the EX laser is locked to the arm cavity length using the AUX PDH servo. The fluctuations in the beatnote between the two lasers is what is monitored as a diagnostic. See Attachment #1. The reference traces in the top pane are from a known good time. The large excess noise between ~80-200 Hz is what I'm concerned about.

A separate issue that can improve the noise is to track down the noise in the 20-80 Hz band - probably some IMC frequency noise issue.

Noise budget:

See Attachment #2

  • I am pretty confident the electronics after the beat mouth are not to blame - I injected a 50 MHz signal from a Marconi and adjusted the signal amplitude to mimic what we get from the beat mouth. The trace labelled "DFD electronics noise" is the noise in this config.
  • The unsuppressed AUX frequency noise was measured with an SR785 (converted to freqnecy noise units knowing the PDH horn-to-horn voltage and the cavity linewidth). I didn't confirm the sensing noise level (dark noise of the AUX PDH loop), but I figure that at 100 Hz (voltage noise of ~100 uV/rtHz on the SR785), we are above the sensing noise level, and so are truly measuring the in-loop frequency noise of the stabilized AUX laser. I also confirmed that the loop UGF was ~10 kHz and phase margin was ~60 degrees, which are nominal numbers.
  • The fact that the excess noise is only in the X arm channel means the PSL frequency is not to blame.

So what could it be? The only things I can think of are (i) the beat mouth photodiode (NF1611) or (ii) excess noise in the fiber carrying the light from EX to the PSL table (but only on this fiber, and not on the EY fiber). Both seem remote to me - I'll test the former by switching the EX and EY fiber inputs to the beat mouth, but apart from this, I'm out of ideas... 

To avoid this kind of issue, we should really have scripted locks of all the basic IFO configs and record the data to summary pages or something - maybe something to do once Guardian is installed, it'd be pretty hacky to do cleanly with shell scripts.

Attachment 1: ALSX_excess.png  94 kB  | Hide | Hide all
ALSX_excess.png
Attachment 2: budget.pdf  47 kB  Uploaded Thu Jan 7 18:56:32 2021  | Hide | Hide all
budget.pdf
ELOG V3.1.3-